1.366×10-1 mol of ammonia (NH3) gas is isolated in a chamber of volume 20.00 mL at 299.1 °C.
Calculate the predicted pressure in the chamber using first the ideal gas law and then the Van der Waals equation.
(For NH3, the Van der Waals constants are: a = 4.17 atm L2 mol-2; b = 0.0371 L mol-1.)
Enter the numerical value (in either decimal or scientific notation) for the ratio, PvdW/Pideal.
1
Expert's answer
2019-10-17T07:18:48-0400

Number of moles of ammonia "(n)=1.366\times 10^{-1}"

Volume of chamber"(V)=20 ml=20\times10^{-3} L"

Temperature"(T)=299.1\degree C=299.1+273=572.1K"

Predicted pressure using ideal gas equation:

"P_{id}V=nRT\ (R=0.0821)"

"P_{id}\times20\times 10^{-3}=1.366\times 10^{-1}\times 0.0821\times 572.1" "\implies P_{id}=313.75\ atm"

Predicted pressure using Van der Wall gas equation:

"(P +\frac{an^2}{V^2})(V-nb)=nRT"

"(P+\frac{4.17\times 0.1366^2}{(20 \times 10^{-3})^2})(20\times 10^{-3}-0.1366\times 0.0371)=" "0.1366 \times 0.0821 \times 572.1"

"P_{van}=235.14\ atm"